# CODI: Compressing Chain-of-Thought into Continuous Space via Self-Distillation

### v1

@@ 4 Experiments
We demonstrate the effectiveness of CODI’s reasoning in a continuous space through experiments on mathematical reasoning tasks.
@@ 4.1 Experimental Setup
We utilize two datasets to train our models–GSM8k-Aug and GSM8k-Aug-NL. (1) We use the GSM8k-Aug dataset from Deng et al. (2023), which has proven effective for training implicit CoT methods Deng et al. (2024); Hao et al. (2024). This dataset extends the original GSM8k training set Cobbe et al. (2021) to 385k samples by prompting GPT-4. To facilitate implicit CoT training, all natural language interleaving within the CoT is removed, leaving only structured mathematical expressions such as “<<10÷5=2>><<2×2=4>><<6×4=24>><<10$\div$ 5=2>><<2$\times$ 2=4>><<6$\times$ 4=24>>< < 10 ÷ 5 = 2 > > < < 2 × 2 = 4 > > < < 6 × 4 = 24 > >”. (2) We also use GSM8k-Aug-NL, a version that preserves natural language explanations, to assess both the generalizability and effectiveness of our approach to compress more verbose CoTs. Examples and statistics are in Appendix C.
@@
In addition to the test split of GSM8k, we assess model robustness on three out-of-domain (OOD) benchmarks: (1) SVAMP Patel et al. (2021), a dataset of elementary-school arithmetic word problems with simple variations designed for robustness test; (2) GSM-HARD Gao et al. (2022), a modified version of the GSM8k test split where numbers are replaced with values of larger magnitude to increase difficulty; and (3) MultiArith Roy and Roth (2015), a subset of MAWPS Koncel-Kedziorski et al. (2016) containing multi-step mathematical word problems. Examples and statistics are in Appendix C.
@@
We consider the following baselines: (1) CoT-SFT: Finetunes the model on CoT data, enabling it to generate intermediate steps followed by the final answer. As CoT-SFT relies on sampling, we set the temperature to 0.1 and report the average result over 10 runs. (2) No-CoT-SFT: Finetunes the model using only direct answers, without generating intermediate steps. (3) iCoT Deng et al. (2024): Implements a curriculum learning strategy called "Stepwise Internalization", which injects CoT’s reasoning patterns into the model’s internal states. This allows the model to generate direct answers with higher accuracy during inference. (4) Coconut Hao et al. (2024): Build upon iCoT by autoregressively generating intermediate continuous CoT representations, similar to the approach in our work. (5) CODI: our method trained with six continuous thought tokens, matching the setup in Coconut. Baselines (2)–(5) are deterministic models, and their results are reported from a single run. Two base models are considered GPT-2 Radford et al. (2019) and LLaMA3.2-1b Meta (2024). More implementation details are in Appendix A.
@@ 4.2 Main Results
Table 1 shows the evaluation results on GSM8k. CODI achieves a significant performance improvement over other implicit CoT methods. In the settings of GPT-2, CODI surpasses iCoT by 45.7% and Coconut by 28.2%. Notably, CODI is the first continuous CoT method to perform on par with CoT-SFT, achieving 99.1% of CoT-SFT’s performance. Unlike iCoT and Cococnut failing to scale up to larger models Hao et al. (2024), CODI successfully scales to LLaMA1b, achieving 90.3% of CoT-SFT’s performance. These results highlight CODI’s superiority in terms of accuracy for in-domain mathematical reasoning tasks.
@@
CODI utilizes a fixed set of six continuous thoughts, enclosed by two special tokens, resulting in a total of eight "tokens" for reasoning. As shown in Table 2, CODI achieves substantial efficiency gains, with a speedup of approximately 2.7× (3.1× CoT compression) for compact CoTs trained on GSM8k-Aug and 5.9× (7.8× CoT compression) for verbose CoTs trained on GSM8k-Aug-NL, demonstrating CODI’s effectiveness in reducing reasoning overhead.
Robustness.
To assess robustness, we evaluate CODI on out-of-distribution datasets. Notably, CODI consistently outperforms CoT-SFT across all three benchmarks for GPT-2. We attribute this to CODI’s reduced tendency to overfit, as evidenced by its significantly lower training accuracy compared to CoT-SFT (Table A3). This difference arises because CODI lacks exact imitation targets for continuous thoughts during training, making it less prone to memorizing patterns and more adaptable to novel scenarios.
Compress CoTs with Natural Language.
Previous works Deng et al. (2024); Hao et al. (2024) primarily trained on GSM8k-Aug, which consists only of mathematical expressions. To evaluate CODI’s generalizability, we extend our analysis to a more complex CoT dataset, GSM8k-Aug-NL. Table 1 shows that both GPT-2 and LLaMA1b perform worse on it compared to GSM8k-Aug. This decrease in performance stems from the additional natural language tokens, which add noise and make imitation learning more difficult. Surprisingly, CODI surpasses CoT-SFT when using GPT-2 and achieves a higher relative score improvement on LLaMA1b compared to models trained on GSM8k-Aug. Moreover, iCoT almost fails in this task because the longer sequence makes curriculum learning challenging. Furthermore, with the average CoT length increasing to 62.1 (Table 2), CODI achieves a compression ratio of 7.8, suggesting that the optimal compression ratio is dataset-dependent. These results demonstrates CODI’s ability to handle more complex CoT training data, showcasing its applicability to diverse reasoning datasets.
Method GSM8k-Aug GSM8k-Aug-NL
Time (#Tokens) Time (#Tokens)
@@
CoT-SFT 0.73s (25.4) 1.62s (68.8)
No-CoT-SFT 0.16s (0) 0.16s (0)
CODI 0.27s (8) 0.27s (8)
Table 2: Efficiency comparison of different reasoning methods in terms of inference time per math problem on GSM8k. Measured with batch size = 1 on an Nvidia A100 GPU. CoT Token counts are shown in parentheses.
Compression Ratio.
The number of continuous thoughts used during training is a crucial hyperparameter, influencing both the computation allocation and the compression ratio. As shown in Figure 3, CODI consistently outperforms Coconut across all compression ratios. Interestingly, both methods exhibit a similar trend: accuracy peaks when using six continuous thoughts. We attribute this to the dataset’s structure, specifically the average number of CoT steps. When fewer than six continuous thoughts are used, the model lacks sufficient expressiveness to capture reasoning steps effectively. Conversely, beyond six, the additional complexity may not provide further benefits, as most problems do not require additional reasoning steps. Instead, the increased sequence length introduces optimization challenges, outweighing any potential gains.
@@ 4.3 Ablation Studies
Table 3: Ablation studies. ind. static teacher refers to training an independent teacher model beforehand. w/ multitask student extends it by allowing the student model to also learn CoT generation. ind. trained teacher refers to training an independent teacher model along with the student model.
@@
To evaluate the need of self-distillation, we tested settings where the student does not share the model with the teacher (Table 3). Without learning explicit CoT generation (ind. static teacher), the model performs badly and fails to generate meaningful continuous CoTs after decoding. Adding an explicit CoT generation objective (w/ multitask student) significantly restores performance, indicating the importance of reference learning. Additionally, training the teacher alongside the student (ind. trained teacher) leads to better results than using a pre-trained, static teacher (ind. static teacher), supporting the argument of the warm-up effect. Finally, using a unified model (CODI) outperforms maintaining separate teacher-student models (ind. trained teacher), reinforcing the idea that shared model representations help mitigate alignment issues in hidden states.
@@
Table 3 shows that removing the L1 loss (Equation 4) linking the teacher and student processes (w/o L1 Loss) leads to a significant performance drop, indicating the importance of supervision from the distillation token. While the model still performs well in CoT generation, it fails to integrate this skill into continuous CoT reasoning, treating them as independent tasks rather than a unified reasoning process.
@@
Keeping the final step of the CoT chain appears to negatively impact performance, supporting our claim that it provides shortcuts. Furthermore, the projection layer of continuous thought tokens (shown as the MLP layer before each of the continous thought token $z_{1}\cdots z_{k}$ in Figure 2) enhances CODI’s effectiveness, likely by helping to discriminate discrete and continuous CoT representations.
5 Further Analysis
Figure 4: A case study illustrating CODI’s interpretability by analyzing its attended tokens and decoded tokens of each of the six latent thought tokens, $z_{1}\cdots z_{6}$. Attended tokens: these represent the top-10 tokens that the continuous thought attends to when generating the next thought/token. Some attended tokens appear in the form of ‘$z_{i}=x$’, indicating attention to the $i$-th continuous thought. Here $x$ represents the top-1 token that the latent thought maps to in vocabulary space. Decoded tokens: these are the top-5 words that the continuous thoughts are projected back to in vocabulary space by multiplying them with the vocabulary embeddings.
5.1 Interpretability Analysis
Interpreting CODI’s continuous thoughts is inherently challenging because these representations lack explicit imitation targets. However, CODI exhibits an ability to produce observable intermediate results (Figure 4) within its continuous thoughts by projecting its last hidden state into vocabulary space via the model’s word embeddings – treating it in the same way as a standard text token. Additionally, the corresponding operands contributing to these intermediate results can often be found in the attended tokens of the latent representation. For example, the second thought token, $z_{2}$, attends to both "1" and "7" to produce the decoded token "7". While the operator itself (e.g., $\times$) is not explicitly visible in the attention mechanism—since operators are in the context—it is reasonable to infer that the transformer layers implicitly perform this operation. Another interesting observation is that each intermediate result is separated by a seemingly meaningless continuous token. We hypothesize that these tokens act as placeholders or transitional states during the computation of intermediate results. This aligns with the idea that the transformer may require multiple passes to complete the calculation for each intermediate step. More case studies are in the Appendix E.
@@ 5.1 Interpretability Analysis
5.2 CODI’s Pattern Learning
GPT-2 CODI Coconut Res Op-Res
Accuracy 43.7% 34.1% 34.0% 35.7%
Table 5: Comparison of GPT-2 finetuned on two datasets derived from CODI’s decoded thoughts. Res: using intermediate results as CoT. Op-Res: using intermediate operators and results as CoT.
Given that CODI’s continuous thoughts can often be decoded into intermediate results, it raises a question: is CODI effectively equivalent to a GPT-2 fine-tuned on a dataset containing CODI’s decoded patterns? We created a dataset containing only intermediate results (e.g., “CoT: 20, 7, 27. Result: 9” translated from the case study in Figure 4). Additionally, since some cases of CODI show decoded operators like ‘$\times$’ and ‘--’ interleaved with intermediate results, we also create a synthetic CoT dataset that includes both operators and results (e.g., “CoT: $\times$, 20, $\times$, 7, $+$, 27. Result: 9”). As shown in Table 5, while models trained on the two synthetic datasets outperform the No-CoT baseline, they give significantly inferior results compared to CoT and CODI, though perform on par with Coconut. These result suggest that CODI learns richer information from the teacher task through distillation than pure imitation on language-level intermediate results alone, highlighting the advantages of our training framework.
@@ 6 Conclusion
We introduced CODI, a novel paradigm for reasoning in the continuous space. Our extensive experiments demonstrate CODI’s effectiveness as the first continuous CoT method to match the performance of explicit CoT, while achieving a high compression ratio. Furthermore, CODI shows its scalability, robustness, generalizability to more complex CoT data, while retaining interpretability. Future research should explore CODI’s application to more diverse and challenging tasks. A promising direction is the integration of multimodality, leveraging continuous embeddings for seamless modality merging. We hope this work inspires further exploration into reasoning in representations more compact than language, paving the way for more efficient and versatile reasoning paradigms.
@@ 7 Limitations
In the paper, our approach focuses on knowledge transfer by probing the token (”:”) responsible for generating the first answer token. However, this choice may be suboptimal, as some answers begin with “-”, and removing such cases improves performance, suggesting that critical reasoning information might also reside in the token generating the second answer token. Additionally, probing the token that concludes the CoT reasoning—potentially summarizing the entire process—could offer alternative supervision signals. Furthermore, the current answer prompt, “The answer is:”, is an arbitrary design choice that may influence the effectiveness of knowledge transfer. Investigating these aspects further could enable CODI to extend its distillation framework to broader reasoning tasks.
@@
Furthermore, CODI’s current configuration is fully deterministic, whereas one advantage of CoT is its inherent stochasticity, which allows models to improve performance through self-consistency, tree search, and Monte Carlo Tree Search (MCTS). Introducing controlled randomness into CODI could restore this property. Potential solutions include integrating dropout layers, leveraging Variational Autoencoders (VAEs), or applying diffusion models atop the current projection layer to enable diverse and robust reasoning paths.
Finally, our studies primarily focus on mathematical problems, as mathematical CoT training data are more abundant and mathematical problems inherently require complex reasoning. However, this focus may limit the generalizability of CODI to broader reasoning scenarios. Future work should explore its applicability to more diverse reasoning tasks beyond mathematical domains.

### v2

@@ 4 Experiments
We demonstrate CODI’s effectiveness in continuous space reasoning through experiments on mathematical and commonsense reasoning tasks.
@@ 4.1 Experimental Setup
We utilize three datasets to train our models–GSM8k-Aug, GSM8k-Aug-NL, and CommonsenseQA-CoT. (1) We use the GSM8k-Aug dataset from Deng et al. (2023), which has proven effective for training implicit CoT methods Deng et al. (2024); Hao et al. (2024). This dataset extends the original GSM8k training set Cobbe et al. (2021) to 385k samples by prompting GPT-4. To facilitate implicit CoT training, all natural language interleaving within the CoT is removed, leaving only structured mathematical expressions such as “<<10÷5=2>><<2×2=4>><<6×4=24>><<10$\div$ 5=2>><<2$\times$ 2=4>><<6$\times$ 4=24>>< < 10 ÷ 5 = 2 > > < < 2 × 2 = 4 > > < < 6 × 4 = 24 > >”. (2) We also use GSM8k-Aug-NL, a version that preserves natural language explanations, to assess both the generalizability and effectiveness of our approach to compress more verbose CoTs. (3) CommonsenseQA-CoT is derived from CommonsenseQA Talmor et al. (2019), a multiple-choice QA dataset built from ConceptNet-based questions Speer et al. (2017). As it lacks CoT annotations, we generate 8.1k CoT examples using GPT-4o-mini, filtered by correctness. The 1.2k-example validation set is used for evaluation. Examples and statistics are in Appendix C.
@@
For mathematical reasoning, we assess model robustness on three out-of-domain (OOD) benchmarks: (1) SVAMP Patel et al. (2021), a dataset of grade-school arithmetic word problems with simple variations designed for robustness test; (2) GSM-HARD Gao et al. (2022), a modified version of the GSM8k test split where numbers are replaced with values of larger magnitude to increase difficulty; and (3) MultiArith Roy and Roth (2015), a subset of MAWPS Koncel-Kedziorski et al. (2016) containing multi-step mathematical word problems. Examples and statistics are in Appendix C.
@@
We consider the following baselines: (1) CoT-SFT: Finetunes the model on CoT data, enabling it to generate intermediate steps followed by the final answer. (2) No-CoT-SFT: Finetunes the model using only direct answers, without generating intermediate steps. (3) iCoT Deng et al. (2024): Implements a curriculum learning strategy called "Stepwise Internalization", which injects CoT’s reasoning patterns into the model’s internal states. This allows the model to generate direct answers with higher accuracy during inference. (4) Coconut Hao et al. (2024): Build upon iCoT by autoregressively generating intermediate continuous CoT representations, similar to the approach in our work. (5) CODI: our method trained with six continuous thought tokens, matching the setup in Coconut. Baseline (1) is sampled 10 times and their average is reported (temperature=0.1), while baselines (2)–(5) are deterministic models, and their results are reported from a single run. Two base models are considered: GPT-2 Radford et al. (2019) and LLaMA3.2-1b-Instruct Meta (2024). More implementation details are in Appendix A.
@@ 4.2 Main Results
From the results on GSM8k in Figure 3 (leftmost column), we observe that CODI largely outperforms existing implicit CoT methods. With both GPT-2 and LLaMA-1b, CODI surpasses Coconut by over 20%. Remarkably, CODI is the first continuous CoT method to achieve performance comparable to CoT-SFT when using GPT-2, reaching 99% of its accuracy. In contrast to iCoT, which fails to scale effectively to larger models, CODI successfully extends to LLaMA-1b, achieving 90% of CoT-SFT performance. These results verify CODI’s effectiveness on in-domain mathematical reasoning tasks.
Compress More Verbose CoTs.
Previous works Deng et al. (2024); Hao et al. (2024) primarily trained on GSM8k-Aug, which consists only of mathematical expressions. To evaluate CODI’s generalizability, we extend our analysis to a more complex CoT dataset, GSM8k-Aug-NL. Figure 3 (2nd column) shows that both GPT-2 and LLaMA-1b perform worse on it compared to GSM8k-Aug. This decrease in performance stems from the additional natural language tokens, which add noise and make imitation learning more difficult. Surprisingly, CODI surpasses CoT-SFT when using GPT-2 and achieves a higher relative score improvement on LLaMA1b compared to models trained on GSM8k-Aug. Moreover, CODI surpasses all other implicit CoT methods, especially at the size of LLaMA-1b, suggesting the effectiveness of self-distillation. Furthermore, with the average CoT length increased to 65.5 (Figure 4), CODI achieves a compression ratio of 8.2, suggesting that the optimal compression ratio is dataset-dependent. These results demonstrate CODI’s ability to handle more complex CoT training data, showcasing its applicability to diverse reasoning datasets.
Commonsense Reasoning.
As shown in Figure 3 (rightmost column), CoT-SFT largely outperforms No-CoT-SFT for GPT-2, which performs nearly random guessing (five choices per question). This indicates that training on CoT benefits GPT-2. Interestingly, CODI surpasses even CoT-SFT. We attribute this to GPT-2’s limited capacity for generating coherent natural language CoTs—CoT-SFT struggles to replicate the quality of the training CoTs, whereas CODI faces less burden by reasoning in a continuous space with fewer tokens. For LLaMA-1b, we observe that CoT data actually hurts performance. We think it is because we force the model to reason in GPT-4o-mini’s pattern which may diverge from LLaMA’s original pattern. Interestingly, CODI outperforms CoT-SFT by a large margin and achieves accuracy comparable to No-CoT-SFT. This shows that our latent reasoning model could better capture intermediate thought processes in continuous spaces, demonstrating the benefit of learning latent representations rather than overfitting of specific CoT patterns.
@@
CODI utilizes a fixed set of six continuous thoughts, enclosed by two special tokens, resulting in a total of eight "tokens" for reasoning. As shown in Figure 4, CODI achieves substantial efficiency gains, with a speedup of approximately 2.7× (3.1× CoT compression) for compact CoTs trained on GSM8k-Aug and 5.9× (8.2× CoT compression) for verbose CoTs trained on GSM8k-Aug-NL, demonstrating CODI’s effectiveness in reducing reasoning overhead.
Figure 4: Efficiency comparison of different reasoning methods in terms of inference time per math problem on GSM8k. Measured with batch size = 1 on an Nvidia A100 GPU. CoT Token counts are shown in parentheses.
Compression Ratio.
The number of continuous thoughts used during training is a crucial hyperparameter, affecting both the computation allocation and the compression ratio. As shown in Figure 5, CODI consistently outperforms Coconut across all compression ratios. Interestingly, both methods exhibit a similar trend: accuracy peaks when using six continuous thoughts. We attribute this to the dataset’s structure, specifically the average number of CoT steps. When fewer than six continuous thoughts are used, the model lacks sufficient expressiveness to capture reasoning steps effectively. Conversely, beyond six, the additional complexity may not provide further benefits, as most problems do not require additional reasoning steps. Instead, the increased sequence length introduces optimization challenges, outweighing any potential gains.
4.3 Out-of-Distribution (OOD) Evaluation
To assess robustness, we evaluate CODI—trained on GSM8k-Aug—on OOD datasets. Remarkably, CODI consistently outperforms all the other implicit CoT baselines and even CoT-SFT across all three OOD benchmarks with GPT-2 (Table 1). Using LLaMA-1b, CODI also performs better compared to iCoT and Coconut. It also demonstrates stronger performance relative to its in-domain results. We attribute CODI’s robustness to its reduced tendency to overfit. Unlike CoT-SFT, which is trained to mimic exact natural language CoT annotations, CODI generates continuous thoughts without direct imitation targets. This lack of rigid supervision likely prevents memorization and promotes greater adaptability to unfamiliar inputs.
Models SVAMP GSM-Hard MultiA
@@ 4.3 Out-of-Distribution (OOD) Evaluation
No-CoT-SFT 44.1 7.1 70.9
CoT-SFT 66.7 15.6 99.3
iCoT 40.9 4.4 39.0
Coconut 48.8 9.9 90.1
CODI 61.1 12.8 96.1
Table 1: Performance comparison (accuracy %) on OOD datasets, i.e., trained on GSM8k-Aug and evaluated on other datasets. The best results are in bold, and the second-best results are underlined.
@@ 4.4 Ablation Studies
Table 2: Ablation studies. ind. static teacher refers to introducing an independently trained teacher model. w/ multitask student allows the student model to also learn CoT generation.
@@
To evaluate the need of self-distillation, we tested settings where the student does not share the model with the teacher (Table 2). Without learning explicit CoT generation (ind. static teacher), the model performs badly and fails to generate meaningful continuous CoTs after decoding. Adding an explicit CoT generation objective (w/ multitask student) significantly restores performance, indicating the importance of reference learning.
@@
Table 2 shows that removing the L1 loss (Equation 5) linking the teacher and student tasks (w/o L1 Loss) leads to a significant performance drop, indicating the importance of supervision from distillation. While the model performs well in CoT generation due to multitask learning, it fails to integrate this skill into continuous CoT reasoning, treating them as independent tasks rather than a unified reasoning process.
@@
Keeping the final step of the CoT chain appears to negatively impact performance, supporting our claim that it provides shortcuts. Furthermore, the projection layer of continuous thought tokens slightly enhances CODI’s effectiveness.
5 Interpretability Analysis
Figure 6: A case study illustrating CODI’s interpretability by analyzing its attended tokens and decoded tokens of each of the six latent thought tokens, $z_{1}\cdots z_{6}$. Attended tokens: these represent the top-10 tokens that the continuous thought attends to when generating the next thought/token. Some attended tokens appear in the form of ‘$z_{i}=x$’, indicating attention to the $i$-th continuous thought. Here $x$ represents the top-1 token that the latent thought maps to in vocabulary space. The model always attends to the first token in the sentence, so we remove that for better visualization. Decoded tokens: these are the top-5 words that the continuous thoughts are projected back to in vocabulary space by multiplying them with the vocabulary embeddings.
Interpreting CODI’s continuous thoughts is inherently challenging because these representations lack explicit imitation targets. However, CODI exhibits an ability to produce observable intermediate results (Figure 6) within its continuous thoughts by projecting its last hidden state into vocabulary space via the model’s word embeddings – treating it in the same way as a standard text token. Additionally, the corresponding operands contributing to these intermediate results can often among the top-ranked attended tokens of the latent representation. For example, the second thought token, $z_{2}$, attends to both "1" and "7" to produce the decoded token "7". While the operator itself (e.g., $\times$) is not explicitly visible in the attention mechanism—since operators are in the context—it is reasonable to infer that the transformer layers implicitly perform this operation. Another interesting observation is that each intermediate result is separated by a seemingly meaningless continuous token. We hypothesize that these tokens act as placeholders or transitional states during the computation of intermediate results. This aligns with the idea that the transformer may require multiple passes to complete the calculation for each intermediate step. More case studies are in the Appendix E.
@@ 6 Conclusion
We introduced CODI, a novel paradigm for reasoning in continuous space. Our extensive experiments demonstrate CODI’s effectiveness as the new SOTA implicit CoT approach, while achieving a high compression ratio. Furthermore, CODI shows its robustness, generalisable to complex datasets, and interpretability. Future research should explore CODI’s application to more diverse and challenging tasks. A promising direction is the integration of multimodality, leveraging continuous representations for seamless modality merging. We hope this work inspires further exploration into reasoning in representations more compact and robust than language, paving the way for more efficient and versatile reasoning paradigms.
@@ 7 Limitations
Implicit CoT methods inherently trade off interpretability compared to explicit CoT. While CODI provides a straightforward probing mechanism for inspecting continuous thoughts, it operates at the token level and faces limitations in reconstructing multi-token entities. For instance, a rare number like 35649 may span multiple tokens due to the tokenizer’s behavior, but the current probing technique only decodes the first token, leaving the remaining components unobserved. More sophisticated probing techniques may be necessary to recover and visualize full semantic units.
Moreover, our approach focuses on knowledge transfer by probing the token (“:”) responsible for generating the first answer token. However, this choice may be suboptimal, as some answers begin with “-”, and removing such cases improves performance, suggesting that critical reasoning information might also reside in the token generating the second answer token. Additionally, probing the token that concludes the CoT reasoning—potentially summarizing the entire process—could offer alternative supervision signals. Furthermore, the current answer prompt, “The answer is:”, is an arbitrary design choice that may influence the effectiveness of knowledge transfer. Investigating these aspects further could enable CODI to extend its distillation framework to broader reasoning tasks.
@@
Finally, while we don’t have sufficient computation resources to scale the training of CODI on larger models, a concurrent paper Geiping et al. (2025b) has demonstrated the feasibility of scaling a latent reasoning model to 3.5B parameters and 800 billion tokens with 4096 GPUs. The resulting model appears to be learning meta-strategies and abstractions for problem solving, as opposed to memorising as in existing LLMs trained on explicit CoT data. This is particularly encouraging, since not all reasoning steps can be easily verbalised (such as visual-spatial reasoning, emotional and social reasoning, and motor reasoning). While Geiping et al. (2025b) focuses on pre-training, we proposed an efficient fine-tuning approach for equipping existing pre-trained LLMs with latent reasoning capabilities.